# Symplectic mapping class relations from pencil pairs

### v1

@@ 3. Symplectic mapping class relations from pencil pairs
In this section we define pencil pairs and then appeal to Theorem 2.4.2 to show how they determine positive relations. Then we will describe some low-dimensional examples.
@@ 3.1. Pencil pairs
Let $X^{\pm}$ be closed varieties with ample line bundles $\operatorname{\mathcal{L}}^{\pm}$ equipped with Lefschetz pencils $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$ of $\operatorname{\mathcal{L}}^{\pm}$ and associated subvarieties $B^{\pm}\subset Z^{\pm}_{i}\subset X^{\pm}$. If there is a symplectomorphism
@@
then we say that $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
Our notation indicates that the pencil pair depends only on the $\operatorname{\mathcal{L}}^{\pm}$ rather than the specified sections $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$, which follows from Lemma 2.4.1.
@@
$(F,\omega_{F})=(Z^{+}_{1}\setminus B,\omega^{+}|_{TZ^{+}})$
as follows: Write $\operatorname{Critv}^{\pm}=\operatorname{Critv}\pi^{\pm}$ for the critical values of our holomorphic functions. We have a collection $\operatorname{\mathcal{L}}^{+}_{i}$ of Lagrangian spheres in $(F,\omega_{F})$ determined by the holomorphic function $\pi^{+}=\boldsymbol{\mathfrak{s}}^{+}_{1}/\boldsymbol{\mathfrak{s}}^{+}_{0}$ and a choice of matching paths for $\operatorname{Critv}^{+}$. We also have a collection $\phi^{-1}\operatorname{\mathcal{L}}^{-}_{i}$ in $(F,\omega_{F})$ determined by $\pi^{-}$ and a choice of matching paths for $\operatorname{Critv}^{-}$. Then appealing to Theorem 2.4.2, we have
@@
In other words, $\vec{L}^{+}$ and $\phi^{-1}\vec{L}^{-}$ gives positive relation. The difference in the lengths of these factorizations is easily counted from Theorem 2.4.2 as
$|\vec{L}^{+}|-|\vec{L}^{-}|=\#\operatorname{Crit}^{+}-\#\operatorname{Crit}^{-}=(-1)^{\dim_{\mathbb{C}}X^{\pm}}(\chi(X^{+})-\chi(X^{-})).$
The formula tells us that whenever we have a pencil pair with $\chi(X^{+})\neq\chi(X^{-})$, then their associated mapping class relations are non-braid-like.
@@ 3.2. The case $\dim_{\mathbb{C}}X^{\pm}=2$
Suppose that $X$ is a closed variety of $\dim_{\mathbb{C}}X=2$ with an ample line bundle $\operatorname{\mathcal{L}}$ having a pencil with associated holomorphic map $\pi$. Then our $Z_{1}$ is a Riemann surface and $B\subset Z_{1}$ is a set of points. Applying Poincaré duality and the adjunction formula,
#$\begin{gathered}\#B=[Z_{1}]^{2}=\int_{Z}\omega_{\operatorname{\mathcal{L}}}=\int_{X}c_{1}(\operatorname{\mathcal{L}})^{2},\quad\chi(Z_{1})=\int_{Z}c_{1}(X)-[Z_{1}]^{2}=\int_{X}(c_{1}(X)-c_{1}(\operatorname{\mathcal{L}}))\wedge c_{1}(\operatorname{\mathcal{L}})\\ \implies\#\operatorname{Crit}\pi=\int_{X}c_{2}(X)+(c_{1}(\operatorname{\mathcal{L}})-2c_{1}(X))\wedge c_{1}(\operatorname{\mathcal{L}}).\end{gathered}$
@@
Theorem 3.2.1.
Let $X^{\pm}$ be a pair of $\dim_{\mathbb{C}}=2$ varieties with ample line bundles $\operatorname{\mathcal{L}}^{\pm}$. Then there is a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair iff the quantities
@@
are independent of the $\pm$ signs.
Using the fact that $c_{1}(\operatorname{\mathcal{L}}^{\otimes}k)=kc_{1}(\operatorname{\mathcal{L}})$, we obtain the following easy lemma.
Theorem 3.2.2.
@@ 3.3. Generalized lantern relations
The following family of example is well-known: Consider $X^{+}=\mathbb{P}^{2},c_{1}(\operatorname{\mathcal{L}}^{+})=2\omega_{\mathbb{P}^{2}}$ and $X^{-}=\mathbb{P}^{1}\times\mathbb{P}^{1},c_{1}(\operatorname{\mathcal{L}}^{-})=(2\pi_{1}^{\ast}+\pi_{2}^{\ast})\omega_{\mathbb{P}^{1}})$. A quick calculation shows that these form a pencil pair by Theorem 3.2.1. The $Z^{\pm}_{1}$ are $\mathbb{P}^{1}$s each having self-intersection number $4$. The contact $3$-manifold given by the boundary of their tubular neighborhoods of the $Z^{\pm}_{1}\subset X^{\pm}$ is the standard lens space $(L(4,1),\xi_{std})$ given by reducing the standard contact $3$-sphere $(S^{3},\xi_{std})=\partial(\mathbb{D}^{4},\beta_{0})$ by the $\mathbb{Z}/4\mathbb{Z}$ action generated by $(z\in\mathbb{S}^{3})\mapsto iz$. The $X^{\pm}\setminus Z^{\pm}_{1}=W^{\pm}$ are its unique symplectic fillings (modulo symplectic blow-up) according to McDuff [M90]. The Lefschetz fibration on $W^{+}$ determined by a holomorphic Lefschetz pencil has fiber a $4$-punctured sphere with three critical points while the Lefschetz fibration on $W^{-}$ is a $4$-punctured sphere with four critical points. As observed by Auroux and Smith [AS08, §5.2] the Lefschetz fibration on $W^{+}$ has monodromy given the left hand side of the lantern relation as it appears in Figure 1. An analysis similar to that of [AS08] reveals that the monodromy of $W^{-}$ is given by the right-hand side of the lantern relation as it appears in Figure 1. From Theorem 3.2.2 we get pencil pairs from the $(X^{+},(\operatorname{\mathcal{L}}^{+})^{\otimes k})$ and $(X^{-},(\operatorname{\mathcal{L}}^{-})^{\otimes k})$ for all $k\in\mathbb{Z}_{>0}$.
Now we describe how the above $k\in\mathbb{Z}_{>0}$ family fits into a larger family of examples, starting with some generalities. Let $\operatorname{\mathcal{L}}_{\Sigma}$ be a line bundle of degree $-d=-d(\operatorname{\mathcal{L}}_{\Sigma})<-2g_{\Sigma}-1$ over a closed Riemann surface $\Sigma$. The $\operatorname{rank}=2$ bundle $\operatorname{Hom}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})=(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})^{\ast}\simeq\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma}^{\ast}$ over the total space of $\mathbb{C}\otimes\operatorname{\mathcal{L}}_{\Sigma}$ descends to a line bundle $\operatorname{\mathcal{L}}$ over the projectivization $\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})$ which is a ruled complex surface $\mathbb{P}^{1}\rightarrow\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})\rightarrow\Sigma$. Sections $\boldsymbol{\mathfrak{s}}$ of $\operatorname{\mathcal{L}}$ take the form $\boldsymbol{\mathfrak{s}}[z_{0},z_{1}]=az_{0}-\boldsymbol{\mathfrak{s}}_{\operatorname{\mathcal{L}}}(z_{1})$ where $\boldsymbol{\mathfrak{s}}_{\operatorname{\mathcal{L}}}$ is a section of $\operatorname{\mathcal{L}}_{\Sigma}^{\ast}\rightarrow\Sigma$, $z_{0}\in\mathbb{C}$, and $z_{1}\in\operatorname{\mathcal{L}}_{\Sigma}$. The condition on $-d$ implies that $\operatorname{\mathcal{L}}_{\Sigma}^{\ast}$ and so $\operatorname{\mathcal{L}}$ are ample. The zero locus of a $a=0$ section is the union of the section $[1,0]$ with $-d$ fibers of the fibration. Desingularizing such a non-generic section, the zero locus $Z$ of a generic section has the topological type of $\Sigma$. Since the section $[1,0]$ has self-intersection $-d$ and the fibers have self-intersection 00, $[Z]\bullet[Z]=d$. Since $\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})$ is a $2$-sphere bundle with a section, $\chi(\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma}))=2\chi(\Sigma)$, eg. from a Gysin sequence. It follows that a pencil for $(\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma}),\operatorname{\mathcal{L}})$ has fiber $\Sigma$ minus $d$ points with $d$ critical points. It’s not difficult to work out that each singular fiber is a nodal curve given by a union of $\Sigma$ with $d-1$ points removed together with a fiber with one point removed. So we see that the total monodromy is a fibered Dehn twist on $\Sigma$ minus $d$ points. In general a pencil for $\operatorname{\mathcal{L}}^{\otimes k}$ the fiber $F_{k}$ is obtained by removing $k^{2}d$ punctures from a closed surface of $\chi=k(\chi(\Sigma)+d)-k^{2}d$. The Lefschetz fibration $\pi_{k}$ will have $2((1-k)\chi(\Sigma)-d)+3k^{2}d$ critical points.
Now consider $\Sigma_{l}$ a degree $l$ curve in $\mathbb{P}^{2}$ (so that $\chi(\Sigma)=l(3-l)$) with $\operatorname{\mathcal{L}}_{\Sigma}^{\ast}=\mathcal{O}_{k}|_{\Sigma_{k}}$ and consider $(X_{l},\operatorname{\mathcal{L}}_{l})$ as so that $d=l^{2}$. The fiber $F_{l,k}$ of a pencil for $(X_{l},\operatorname{\mathcal{L}}_{l}^{k})$ is then obtained from removing $(kl)^{2}$ points from a closed Riemann surface of $\chi=k(l(3-l)+l^{2})-(kl)^{2}$, with the associated Lefschetz fibration $\pi_{l,k}$ having #$\#\operatorname{Crit}\pi_{l,k}=2((1-k)l(3-l)-l^{2})+3(kl)^{2}$. In the case $l=2$, $(X_{l},\operatorname{\mathcal{L}}_{l})$ is the second Hirzebruch surface, being identical to the $(X^{-},\operatorname{\mathcal{L}}^{-})$ above. So for each such $l\geq 2$, the $(\mathbb{P}^{2},\mathcal{O}_{l})$ and $(X_{l},\operatorname{\mathcal{L}}_{l})$ determine a pencil pair. Applying Theorem 3.2.2, the $(\mathbb{P}^{2},\mathcal{O}_{l}^{\otimes k}=\mathcal{O}_{lk})$ and $(X_{l},\operatorname{\mathcal{L}}_{l}^{\otimes k})$ then determine pencil pairs for all $l\geq 2,k\geq 1$, summarized in the following theorem.
@@
Let $F_{l,k}$ be an oriented surface of genus $g_{l,k}=1+\frac{1}{2}kl(kl-3)$ with $(kl)^{2}$ boundary components. Then a boundary Dehn twist $\tau_{\partial F_{l,k}}$ can be factorized as products of
@@
    $(kl)^{2}$ positive Dehn twists along the individual boundary components of $F_{k,l}$ determined by $\tau_{\partial F_{k,l}}$,
@@
    $3(kl)^{2}-6kl-2l^{2}+6l$ positive Dehn twists determined by a pencil for $(X_{l},\operatorname{\mathcal{L}}_{l}^{k})$, and
@@
    $3((kl)^{2}-3(kl)+1)$ positive Dehn twists determined by a pencil $(\mathbb{P}^{2},\mathcal{O}_{kl})$.
For the $k=1$ case the first two collections of curves coincide, and when $l=2,k=1$ these collections give the classical lantern relation.
Factorizing a given $kl\in\mathbb{Z}_{>1}$ in different ways then gives us many factorizations of $\tau_{\partial F_{l,k}}$. For example when $k=2^{n-j}$ and $l=2^{j}$ for $j=1,\dots,n-1<n$ the above theorem gives us factorizations into
$m=2^{n},3(2^{2n}-2^{n+1})+2^{j+1},3(2^{2n}-3\cdot 2^{n}+1)$
Dehn twists. By looking at the compliments of $\operatorname{\mathcal{L}}^{2^{n-j}}_{2^{j}}$ hyperplane sections in the $X_{l}$ and $\mathcal{O}_{2^{n}}$ hyperplane sections in $\mathbb{P}^{2}$, we see that the boundary of a tubular neighborhood of a $\deg=2^{n}$ curve in $\mathbb{P}^{2}$ has at least $n$ distinct (algebraic) Weinstein fillings, distinguishable by their Euler characteristics.
4. Pencil pairs from projective models
Our technique for finding new pencil pairs in this article will rely on our identification of the $Z^{\pm}_{1}\subset X^{\pm}$ with complete intersections in “big” projective varieties $\mathcal{M}$. In this section we describe this technique and use it show a infinite families of pencil pairs can be determined from a single pencil pair, analogous to the case of Theorem 3.2.2.

### v2

@@ 3. Symplectic mapping class relations from pencil pairs
In this section we define pencil pairs and then appeal to Theorem 2.4.2 to show how they determine positive relations (Definition 2.1.1). Then we will describe some low-dimensional examples.
@@ 3.1. Pencil pairs
Let $X^{\pm}$ be closed varieties equipped with very ample line bundles $\operatorname{\mathcal{L}}^{\pm}$, Lefschetz pencils $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$ for the $\operatorname{\mathcal{L}}^{\pm}$, and associated subvarieties $B^{\pm}\subset Z^{\pm}_{i}\subset X^{\pm}$. If there is a symplectomorphism
@@
we say that $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
It is implicit in our notation that the pencil pair depends only on the $\operatorname{\mathcal{L}}^{\pm}$ rather than the specified sections $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$. This abuse of notation is permitted by Lemma 2.4.1.
@@
$(F,\omega_{F})=\left(Z^{+}_{1}\setminus B,\omega^{+}|_{TZ_{1}^{+}}\right)$
as follows: We have a collection $\vec{L}^{+}=(L^{+}_{i})_{i=1}^{\#\operatorname{Crit}\pi^{+}}$ of Lagrangian spheres in $(F,\omega_{F})$ determined by the holomorphic function $\pi^{+}=\boldsymbol{\mathfrak{s}}^{+}_{1}/\boldsymbol{\mathfrak{s}}^{+}_{0}$ and a choice of matching paths for $\operatorname{Critv}\pi^{+}$. Here we recall that $F$ is a smooth fiber of $\pi^{+}$. We also have a collection $\phi^{-}\vec{L}^{-}=(\phi^{-1}L^{-}_{i})_{i=1}^{\#\operatorname{Crit}\pi^{-}}$ of Lagrangian spheres in $(F,\omega_{F})$ determined by $\pi^{-}$ and a choice of matching paths for $\operatorname{Critv}\pi^{-}$. Then by Theorem 2.4.2,
@@
In other words, the pair $\vec{L}^{+},\phi^{-1}\vec{L}^{-}$ is a positive relation. Theorem 2.4.2 also yields
$|\vec{L}^{+}|-|\vec{L}^{-}|=\#\operatorname{Crit}\pi^{+}-\#\operatorname{Crit}\pi^{-}=(-1)^{\dim_{\mathbb{C}}X^{\pm}}(\chi(X^{+})-\chi(X^{-})).$
In light of the results of §2.2, this formula tells us that whenever we have a pencil pair with $\chi(X^{+})\neq\chi(X^{-})$, then the associated mapping class relation is non-braid-like.
@@ 3.2. The case $\dim_{\mathbb{C}}X^{\pm}=2$
Suppose that $X$ is a closed variety of $\dim_{\mathbb{C}}X=2$ equipped with a very ample line bundle $\operatorname{\mathcal{L}}$ having a pencil with associated holomorphic map $\pi$. Then our divisor $Z_{1}$ is a Riemann surface and base locus $B\subset Z_{1}$ is a set of points. Applying Poincaré duality and the adjunction formula,
(3.2.1) #$\begin{gathered}\#B=[Z_{1}]^{2}=\int_{Z_{1}}\omega_{\operatorname{\mathcal{L}}}=\int_{X}c_{1}(\operatorname{\mathcal{L}})^{2},\\ \chi(Z_{1})=\int_{Z_{1}}c_{1}(X)-[Z_{1}]^{2}=\int_{X}(c_{1}(X)-c_{1}(\operatorname{\mathcal{L}}))\wedge c_{1}(\operatorname{\mathcal{L}})\\ \implies\#\operatorname{Crit}\pi=\int_{X}c_{2}(X)+(c_{1}(\operatorname{\mathcal{L}})-2c_{1}(X))\wedge c_{1}(\operatorname{\mathcal{L}}).\end{gathered}$
@@
Lemma 3.2.1.
Let $X^{\pm}$ be a pair of $\dim_{\mathbb{C}}=2$ varieties equipped with very ample line bundles ±$\operatorname{\mathcal{L}}^{\pm}\to X^{\pm}$. Then there is a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair iff the quantities
@@
are independent of their $\pm$ signs.
Using the fact that $c_{1}(\operatorname{\mathcal{L}}^{\otimes k})=kc_{1}(\operatorname{\mathcal{L}})$, the following lemma follows easily from Equation (3.2.1).
Lemma 3.2.2.
@@ 3.3. Generalized lantern relations
The following family of examples is well-known: Consider
$X^{+}=\mathbb{CP}^{2},\quad c_{1}(\operatorname{\mathcal{L}}^{+})=2\omega_{\mathbb{CP}^{2}},\quad X^{-}=\mathbb{CP}^{1}\times\mathbb{CP}^{1},\quad c_{1}(\operatorname{\mathcal{L}}^{-})=(2\pi_{1}^{\ast}+\pi_{2}^{\ast})\omega_{\mathbb{CP}^{1}}.$
@@
Here the $\pi_{i}$ denote projections onto the $\mathbb{CP}^{1}$ factors of $X^{-}$. A quick calculation shows that these $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ form a pencil pair by Lemma 3.2.1. The $Z^{\pm}_{1}$ are $\mathbb{CP}^{1}$s each having self-intersection number $4$. The contact $3$-manifold given by the boundary of a tubular neighborhood of a $Z^{\pm}_{1}\subset X^{\pm}$ is the standard lens space $(L(4,1),\xi_{std})$ given by reducing the standard contact $3$-sphere $(S^{3},\xi_{std})=\partial(\mathbb{D}^{4},\beta_{0})$ by the $\mathbb{Z}/4\mathbb{Z}$ action generated by $z\mapsto iz$ for $z\in\mathbb{S}^{3}\subset\mathbb{C}^{2}$. The $W^{\pm}=X^{\pm}\setminus Z^{\pm}_{1}$ are its unique symplectic fillings (modulo symplectic blow-up) according to McDuff [M90]. The Lefschetz fibration on $W^{+}$ determined by a holomorphic Lefschetz pencil has fiber a $4$-punctured sphere with three critical points. As noted after Theorem 2.4.2, $W^{+}$ is symplectomorphic to a $\mathbb{D}^{\ast}\mathbb{RP}^{2}$. The Lefschetz fibration on $W^{-}$ has fiber a $4$-punctured sphere with four critical points. As observed by Auroux and Smith [AS08, §5.2] the Lefschetz fibration on $W^{+}$ has monodromy given the left hand side of the lantern relation as it appears in Figure 1. An analysis similar to that of [AS08] reveals that the monodromy of the Lefschetz fibration on $W^{-}$ is given by the right-hand side of the lantern relation as it appears in Figure 1. See [T20, §5]. Lemma 3.2.2 yields pencil pairs for the the $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k})$ for each $k\in\mathbb{Z}_{>0}$. We can view these pencil pairs as generalizing the lantern relation, since for any $k$ the difference in the number of critical points is always $\chi(X^{-})-\chi(X^{+})=1$.
Remark 3.3.1.
As we learned from reading [L97], the lantern relation was first found by Dehn [D87, p.333] and then later rediscovered by Johnson [J79], to whom it is often attributed.
An alternative, higher-dimensional generalization of the lantern relation (associated to the total space of $\mathbb{D}^{\ast}\mathbb{CP}^{2}$ rather than $\mathbb{D}^{\ast}\mathbb{RP}^{2}$) is described in [T20].
Now we describe how the above $k\in\mathbb{Z}_{>0}$ family fits into a larger collection of examples starting with a general construction. Let $\Sigma$ be a closed Riemann surface of Euler characteristic $\chi$, choose some integer $d>1+g_{\Sigma}=3-\chi$, and let $\ell_{\chi,d}\to\Sigma$ be a line bundle of $\deg\ell_{\chi,d}=-d$. Therefore the dual line $\ell_{\chi,d}^{\ast}\to\Sigma$ has $\deg=d$. Write $X_{\chi,d}=\mathbb{P}(\mathbb{C}\oplus\ell_{\chi,d})$ for the projectivization, which is a complex ruled surface,
[…]
See for example [MS99, Example 4.26]. Since $X_{\chi,d}$ is a $2$-sphere bundle with a section it follows, e.g. from a Gysin sequence, that $\chi(X_{\chi,d})=2\chi$. The rank $2$ bundle $(\mathbb{C}\oplus\ell_{\chi,d})^{\ast}\simeq\mathbb{C}\oplus\ell_{\chi,d}^{\ast}$ over the total space of $\mathbb{C}\oplus\ell_{\chi,d}$ descends to a line bundle $\operatorname{\mathcal{L}}_{\chi,d}\to X_{\chi,d}$. Sections $\boldsymbol{\mathfrak{s}}\in H^{0}(X_{\chi,d},\operatorname{\mathcal{L}}_{\chi,d})$ of $\operatorname{\mathcal{L}}_{\chi,d}$ have the form
$\boldsymbol{\mathfrak{s}}[z_{0},z_{1}]=az_{0}-\boldsymbol{\mathfrak{s}}^{\ast}(z_{1}),\quad a,z_{0}\in\mathbb{C},\quad z_{1}\in\ell_{\chi,d},\quad\boldsymbol{\mathfrak{s}}^{\ast}\in H^{0}(\Sigma,\ell_{\chi,d}^{\ast}).$
Since $\ell_{\chi,d}^{\ast}$ is very ample, so is $\operatorname{\mathcal{L}}_{\chi,d}$. A section $\boldsymbol{\mathfrak{s}}_{si}$ of $\operatorname{\mathcal{L}}_{\chi,d}$ for which $a=0$ and $\boldsymbol{\mathfrak{s}}^{\ast}$ is generic will have singular zero locus $Z_{si}$ whose irreducible components are the $[1:0]$ submanifold together with $d$ fibers of $p$. Take a generic $a\neq 0$ section $\boldsymbol{\mathfrak{s}}_{sm}$ close to $\boldsymbol{\mathfrak{s}}_{si}$ whose zero locus $Z_{sm}$ is smooth. Then $Z_{sm}$ is obtained by smoothing the singularities of $Z_{si}$, so $\chi(Z_{sm})=\chi$ and $\operatorname{PD}_{X_{\chi,d}}[Z_{sm}]=c_{1}(\operatorname{\mathcal{L}}_{\chi,d})$. Analyzing the intersections of $Z_{sm}$ and $Z_{si}$ (which have the same $H_{2}(X_{\chi,d})$ class) yields $\int_{X_{\chi,d}}c_{1}(\operatorname{\mathcal{L}}_{\chi,d})^{2}=d$. The adjunction formula applied to $Z_{sm}$ gives $c_{1}(TX_{\chi,d})=\chi+d$. Combining our calculations with Equation (3.2.1), a Lefschetz fibration associated to a pencil $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ for $\operatorname{\mathcal{L}}_{\chi,d}^{\otimes k}\to X_{\chi,d}$ has
  1. (1)
    $Z_{0,\chi,d,k}=\boldsymbol{\mathfrak{s}}_{0}^{-1}(0)$ with $[Z_{0,\chi,d,k}]=dk^{2}$ and $\chi(Z_{0,\chi,d,k})=k(\chi+d-dk)$,
  2. (2)
    total space $W_{\chi,d,k}=X_{\chi,d}\setminus Z_{0,\chi,d,k}$ with $\chi(W_{\chi,d,k})=2\chi-k\chi-kd+dk^{2}$,
  3. (3)
    #$\#\operatorname{Crit}=3dk^{2}+2(\chi-kd-k\chi)$ critical points of $\pi=\boldsymbol{\mathfrak{s}}_{1}/\boldsymbol{\mathfrak{s}}_{0}:X_{\chi,d,k}\to\mathbb{C}$, and
  4. (4)
    non-singular fiber of $\pi$ of genus $g_{\chi,d,k}=1+\frac{k}{2}(\chi+d-dk)$ with $dk^{2}$ punctures.
Note that $X_{2,2}\to\mathbb{CP}^{1}$ is a trivial fibration, and the symplectic manifold we get is the second Hirzebruch surface, identical to $(X^{-},\omega_{\operatorname{\mathcal{L}}^{-}})$ above. A quick calculation shows that for any $m,k>0$, $(\mathbb{CP}^{2},\mathcal{O}_{mk})$ and $(X_{\chi,m^{2}},\operatorname{\mathcal{L}}_{\chi,m^{2}}^{\otimes k})$, $\chi=m(3-m)$ are a pencil pair via Lemma 3.2.1.
@@
Let $F_{m,k}$ be an oriented surface of genus $g_{m,k}=1+\frac{1}{2}mk(mk-3)$ with $(mk)^{2}$ boundary components. Writing $\chi=m(3-m)$, a boundary Dehn twist $\tau_{\partial F_{m,k}}$ can be factorized as products of
@@
    $(mk)^{2}$ Dehn twists along the individual boundary components of $F_{m,k}$ determined by $\tau_{\partial F_{m,k}}$,
@@
    $3(mk)^{2}-2m(3k+m-3)$ Dehn twists determined by a pencil for $(X_{\chi,m^{2}},\operatorname{\mathcal{L}}_{\chi,m^{2}}^{\otimes k})$, and
@@
    $3((mk)^{2}-2mk+1)$ Dehn twists determined by a pencil $(\mathbb{CP}^{2},\mathcal{O}_{mk})$.
For the $k=1,m=2$ case the first two collections of curves coincide, and the collections of curves from the last two items give the classical lantern relation.
By factoring $mk$ in different ways, the above constructions can give many holomorphic Weinstein fillings $W_{\chi,m^{2},k},\mathbb{CP}^{2}\setminus\Sigma_{mk}$ of the contact $3$-manifold given as the ideal boundary of the complement of a $\deg=mk$ curve $\Sigma_{mk}\subset\mathbb{CP}^{2}$. For example, choose $N\in\mathbb{Z}_{\geq 2}$ and $m=2^{i},k=2^{N-i}$ for $0<i<N$. Then
$\chi(W_{\chi,2^{2i},2^{N-i}})=2^{i+1}(3-2^{i})-2^{N+1},\quad\chi(\mathbb{CP}^{2}\setminus\Sigma_{2^{N}})=3+2^{N}(2^{N}-3).$
These Euler characteristics are all distinct, implying the following theorem.
Theorem 3.3.3.
Let $(M,\xi)$ be the ideal contact boundary of a tubular neighborhood of a $\deg=2^{N}$ curve in $\mathbb{CP}^{2}$ with $N\geq 2$. Then $(M,\xi)$ has at least $N$ distinct (holomorphic) Weinstein fillings.
4. Pencil pairs from projective models and high-degree hypersurfaces
The existence of our $\dim_{\mathbb{C}}X^{\pm}=3$ pencil pairs (determining mapping class relations on Weinstein $4$-manifolds) relies on identifications of divisors $Z^{\pm}_{1}\subset X^{\pm}$ with complete intersections in “big” projective varieties $\mathcal{M}$. In §4.1 we describe this technique. In §4.2 we show that a single pencil pair found using this technique determines an infinite family of pencil pairs, generalizing Lemma 3.2.2 to all dimensions.